### v1

# TurkicNLP: An NLP Toolkit for Turkic Languages

Abstract
Natural language processing for the Turkic language family, spoken by over 200 million people across Eurasia, remains fragmented, with most languages lacking unified tooling and resources. We present TurkicNLP, an open-source Python library providing a single, consistent NLP pipeline for Turkic languages across four script families: Latin, Cyrillic, Perso-Arabic, and Old Turkic Runic. The library covers tokenization, morphological analysis, part-of-speech tagging, dependency parsing, named entity recognition, bidirectional script transliteration, cross-lingual sentence embeddings, and machine translation through one language-agnostic API. A modular multi-backend architecture integrates rule-based finite-state transducers and neural models transparently, with automatic script detection and routing between script variants. Outputs follow the CoNLL-U standard for full interoperability and extension. Code and documentation are hosted at https://github.com/turkic-nlp/turkicnlp.
TurkicNLP: An NLP Toolkit for Turkic Languages
Sherzod Hakimov Computational Linguistics University of Potsdam firstnamelastname@gmail.com
1 Introduction
1.1 Background
The Turkic languages form one of the world’s largest language families, comprising over 35 recognized languages spoken by approximately 200 million people across a vast geographic region stretching from Turkey and the Balkans to Siberia and northwest China Johanson (2021). These languages are characterized by shared typological features such as agglutinative morphology, vowel harmony, and a subject-object-verb (SOV) word order (as illustrated in Table 1), which distinguish them significantly from many Indo-European languages Jurafsky and Martin (2026). Furthermore, the Turkic family exhibits remarkable script diversity, utilizing Latin, Cyrillic, and Perso-Arabic alphabets, with historical use of Old Turkic Runic scripts. This linguistic and scriptural richness presents unique challenges and opportunities for Natural Language Processing (NLP).
Language (ISO) Sentence Language (ISO) Sentence
Latin script
Turkish (tur) Ali kitabı gördü. Azerbaijani (aze) Ali kitabı gördü.
Turkmen (tuk) Ali kitaby gördi. Gagauz (gag) Ali kitabı gördü.
Crimean Tatar (crh) Ali kitanı kördi. Karakalpak (kaa) Ali kitaptı ko’rdi.
Uzbek (uzb) Ali kitobni ko’rdi.
Cyrillic script
Kazakh (kaz) Aли кітапты көрді. Kyrgyz (kir) Али китепти көрдү.
Tatar (tat) Али китапны күрде. Bashkir (bak) Али китапты күрҙе.
Nogai (nog) Али китапты коьрди. Kumyk (kum) Али китапны гёрдю.
Karachay-Balkar (krc) Али китапны кёрдю. Altai (alt) Али бичикти кӧрди.
Sakha (sah) Али кинигэни көрдө. Tuvan (tyv) Али номну көрген.
Chuvash (chv) Али кӗнекене курнӑ. Khakas (kjh) Али кітапты көрді.
Perso-Arabic script (right-to-left)
Uyghur (uig) ئەلى كىتابنى كۆردى.
Table 1: The sentence “Ali saw the book” across 20 Turkic languages, illustrating the SOV word order (subject Ali — object book with accusative suffix — verb saw at the end) and the three active script families: Latin, Cyrillic, and Perso-Arabic.
Figure 1: Geographic distribution of the Turkic language family (Source: Wikimedia Commons).
1.2 Motivation
Despite the size and diversity of the Turkic language family, NLP tools and resources remain heavily skewed toward Turkish, with most other languages severely underresourced Joshi et al. (2020). This fragmentation means that while some languages like Turkish have access to mature NLP toolkits such as Zemberek Akın and Akın (2007), and others like Kazakh benefit from specialized tools developed by institutions like ISSAI, there is no single, unified library that covers the breadth of the Turkic family with a consistent API. Existing multilingual platforms like Apertium offer rule-based machine translation and morphological analyzers for many Turkic languages Forcada et al. (2011), but often lack a direct, user-friendly Python API for pipeline integration. The agglutinative nature of Turkic languages demands specialized morphological analysis approaches that go beyond the tokenization and part-of-speech tagging paradigms common in NLP for less morphologically complex languages. Moreover, ongoing script transitions in countries like Kazakhstan and Uzbekistan necessitate robust and unified transliteration tooling to bridge linguistic data across different orthographies. The absence of a comprehensive, accessible, and unified NLP toolkit for Turkic languages impedes research, resource creation, and the development of practical applications, particularly for the numerous low-resource languages within the family.
1.3 Contributions
This work makes the following contributions:
  • Unified Pipeline API: A single Pipeline interface covering tokenization, morphological analysis, POS tagging, lemmatization, dependency parsing, named entity recognition, transliteration, sentence embeddings, and machine translation for 24 Turkic languages—with the same calling convention regardless of language or backend.
  • Script-Aware Architecture: Automatic script detection and routing across Latin, Cyrillic, Perso-Arabic, and Old Turkic Runic, with bidirectional transliteration for 8 language–script pairs.
  • Multi-Backend Integration: Rule-based Apertium FST morphological analyzers for 20 languages and Stanza neural models for the 5 languages with Universal Dependencies treebanks, selectable without API changes. NLLB-200 provides cross-lingual embeddings and translation for all 24 languages.
  • CoNLL-U Interoperability: A Stanza-compatible hierarchical document model with native CoNLL-U import and export, ensuring compatibility with Universal Dependencies tools and treebanks.
  • Open-Source Release: Apache-2.0 licensed, with documentation and code at https://github.com/turkic-nlp/turkicnlp and companion Jupyter notebooks for all supported languages at https://github.com/turkic-nlp/turkic-nlp-code-samples.
2 Related Work
2.1 General-Purpose NLP Toolkits
General-purpose NLP toolkits have reached high maturity for major European languages but offer limited coverage of the Turkic family. Stanza Qi et al. (2020) provides neural pipeline components trained on Universal Dependencies treebanks for 66+ languages, including five Turkic languages: Turkish, Kazakh, Kyrgyz, Uyghur, and Ottoman Turkish. However, it lacks a morphology-first design, does not support script transliteration, and provides no integration with Apertium FSTs for the remaining Turkic languages. Trankit Nguyen et al. (2021) is a transformer-based multilingual toolkit covering 56 languages via pipeline components built on XLM-R; it overlaps with Stanza on UD-treebank Turkic languages but similarly provides no script conversion or FST morphology. spaCy Honnibal et al. (2020) is an industrial-strength library optimized for efficiency across major European languages but offers no dedicated Turkic NLP support beyond standard Unicode handling. NLTK Bird et al. (2009) is a widely-used educational framework with minimal Turkic-specific components. Polyglot Al-Rfou’ et al. (2013) provides distributed word representations and basic NLP components for 100+ languages, including several Turkic varieties, but does not support morphological analysis, script conversion, or unified pipeline integration. UDPipe Straka et al. (2016) offers trainable pipelines for any CoNLL-U treebank but does not bundle Turkic models and lacks transliteration support. UralicNLP Hämäläinen (2019), built on the Giellatekno/Divvun FST infrastructure, provides a comparable morphology-first, script-aware toolkit for Uralic minority languages; its architecture—FST analyzers accessed through a unified Python API—directly parallels TurkicNLP’s design, though it targets an entirely different language family with different morphological and script profiles. calamanCy Miranda (2023) and PyThaiNLP Phatthiyaphaibun et al. (2023) exemplify a complementary paradigm: language-specific pipelines that consolidate disjoint NLP tools for morphologically complex under-resourced languages behind a unified API. TurkicNLP occupies a similar role—consolidating fragmented tools for under-resourced languages—but extends the paradigm to a 24-language family with integrated multi-backend swappability and script-aware transliteration routing as first-class architectural features.
2.2 Turkic-Specific NLP Tools and Pipelines
Several tools address individual Turkic languages, organized here by language.
Turkish: Zemberek-NLP Akın and Akın (2007) is a comprehensive Java library for Turkish NLP covering morphological analysis, POS tagging, and spell checking; it is restricted to Turkish, does not expose a pip-installable Python API, and has no cross-family support. TurkishDelightNLP Alecakir et al. (2022) is a neural Turkish NLP pipeline offering tokenization, morphological analysis, dependency parsing, and NER with strong performance on Turkish; it targets Turkish exclusively and does not integrate with other Turkic languages or script conversion. VNLP Turker et al. (2023) is an open-source Python NLP library for Turkish providing morphological analysis, named entity recognition, dependency parsing, and sentiment analysis, including quantitative benchmarks on standard Turkish test sets; like TurkishDelightNLP it is Turkish-only and provides no script handling. Istanbul Technical University (ITU) [^1] https://bbf.itu.edu.tr/en/research/research-labs/nlp-lab maintains a complementary Turkish NLP infrastructure including neural pipelines, the IMST and ITU Web treebanks, TRopBank PropBank, and supporting corpora Eryiğit et al. (2023). The Turkish NLP Suite Demir (2023) is an open-source non-profit collection that includes spaCy-compatible Turkish models, the SentiTurca sentiment corpus, TrGLUE benchmark, InstructTurca instruction-tuning data, and BellaTurca.
Kazakh: The Institute of Smart Systems and Artificial Intelligence (ISSAI) at Nazarbayev University [^2] https://issai.nu.edu.kz/ has developed one of the most comprehensive Turkic-language NLP ecosystems. Text resources include KazNERD Yeshpanov et al. (2022) (112,702 tokens, 25 entity classes), KazQAD Yeshpanov et al. (2024a), and KazParC Yeshpanov et al. (2024c) (371,902 parallel sentences across Kazakh, English, Russian, and Turkish). The speech strand covers KazakhTTS2 Mussakhojayeva et al. (2022) (271 hours TTS), a multilingual Turkic speech corpus spanning 10 languages Mussakhojayeva et al. (2023a), and the Söyle noise-robust ASR system for 11 Turkic languages Mussakhojayeva et al. (2023b).
Kyrgyz: KyrgyzNER Turdubaeva et al. (2024) provides 10,900 annotated sentences across 27 entity classes for Kyrgyz NER.
Multiple languages: Apertium Forcada et al. (2011) provides rule-based morphological analyzers and machine translation for 20+ Turkic languages via finite-state transducers, representing the most comprehensive freely available resource for Turkic morphological analysis. Its native interface is command-line-based, loading Apertium FST binaries via the hfst Python package without requiring system-level installation. Abdurakhmonova et al. 2022 present an NLP tool for linguistic analysis of Turkic languages focusing on morphological annotation, though it addresses a limited language set without a Python package interface. Universal Dependencies treebanks for Turkic languages were created by dedicated community efforts: the Turkish IMST treebank Sulubacak et al. (2016), the Tatar NMCTT treebank Taguchi et al. (2022), the Uyghur UDT Eli et al. (2016), the Yakut/Sakha treebank Merzhevich and Ferraz Gerardi (2022), and Old Turkish Derin and Harada (2021), with a parallel aligned treebank for Azerbaijani, Kyrgyz, Turkish, and Uzbek Eslami et al. (2025). These contributions remain language-specific and are not integrated into a unified pipeline toolkit.
2.3 Multilingual Models and Benchmarks
Large-scale multilingual models and language-specific pretrained models now cover a growing number of Turkic languages, organized here by language.
Turkish: BERTurk Schweter (2020) provides multiple Turkish BERT variants (cased, uncased, 128k vocabulary). TURNA Uludoğan et al. (2024) provides a 1.1B-parameter Turkish encoder-decoder language model. Large generative LLMs include Kumru-2B Turker et al. (2024) and Trendyol-LLM Trendyol Tech (2024).
Kazakh: KAZ-LLM Yeshpanov et al. (2024b) (8B and 70B parameters, LLaMA 3.1-based) and Sherkala-Chat Koto et al. (2025) are generative models targeting Kazakh.
Azerbaijani: aLLMA Hajili et al. (2024) is a language model for Azerbaijani.
Uzbek: UzBERT Mansurov and Mansurov (2021), TahrirchiBERT Mamasaidov et al. (2021), and BERTbek Kuriyozov et al. (2024) cover Uzbek.
Kyrgyz: KyrgyzBERT Metinov (2023) is available for Kyrgyz.
Multilingual: Large-scale multilingual models such as mBERT Devlin et al. (2019) and XLM-R Conneau et al. (2020) provide broad cross-lingual representations for Turkic NLP, but suffer from high tokenizer fertility on agglutinative languages and do not constitute processing pipelines. The mGPT multilingual generative model Shliazhko et al. (2024) covers nine Turkic languages (Azerbaijani, Bashkir, Chuvash, Kazakh, Kyrgyz, Tatar, Turkmen, Uzbek, and Sakha). The NLLB-200 model NLLB Team et al. (2022) covers 200 languages including 14 Turkic variants for machine translation.
Evaluation benchmarks: Evaluation benchmarks for Turkic NLP have advanced considerably. TUMLU Isbarov et al. (2025) covers 9 Turkic languages with 38K multiple-choice questions; Kardeş-NLU Senel et al. (2024) assesses cross-lingual NLU transfer across Turkic languages; and TR-MMLU Bayram et al. (2025) provides 6,200 multiple-choice questions across 62 subject categories specifically for benchmarking Turkish LLMs. Veitsman and Hartmann 2025 survey the state of NLP for Central Asian Turkic languages (Kazakh, Kyrgyz, Uzbek, Uyghur, Turkmen), documenting persistent resource gaps and the challenges of cross-lingual transfer from higher-resource Turkic languages. Collectively, these benchmarks underscore the need for unified processing infrastructure to support training and evaluation pipelines.
2.4 Resource Collections and Community Initiatives
Several community-driven collections catalogue the growing body of Turkic NLP resources. The Turkish NLP resource list Turkish NLP Community (2023) [^3] https://github.com/maidis/nlp-tr is a community-maintained index of Turkish NLP tools, corpora, treebanks, and pre-trained models. The Turkic Interlingua (TIL) project Syzdykov and et al. (2021)  [^4] https://github.com/turkic-interlingua/til-mt provides parallel corpora across 22 Turkic languages together with baseline systems for machine translation, automatic speech recognition, and text-to-speech, enabling systematic multilingual research across the family. Organisational and workshop infrastructure complements these datasets: SIGTURK, the ACL Special Interest Group on Turkic NLP SIGTURK Organizing Committee (2026) [^5] https://sigturk.github.io, organises dedicated workshop venues that bring together researchers working on Turkic languages and provide an important venue for disseminating resources and evaluation results.
2.5 Turkic NLP Resource Landscape
Table LABEL:tab:resources-full (Appendix) catalogues datasets, NLP components, and toolkit coverage for all 24 languages, grouped by genetic branch. Table LABEL:tab:llm-models (Appendix) surveys all publicly available pretrained neural language models—encoder-only, generative, and multilingual—for each supported language. Together, these tables motivate the design choices described in the subsequent sections.
3 Library Design and Architecture
3.1 Design Goals
TurkicNLP was designed around six core principles. First, a unified API: a single Pipeline entry point operates identically across all 24 supported languages, abstracting backend differences and enabling users to switch languages without code changes. Second, multi-backend: each processor can be fulfilled by interchangeable backends—rule-based, Apertium FST, or Stanza neural—without modifying user-facing code; backend selection is controlled either via catalog.json defaults or explicit constructor parameters. Third, script-awareness: the library detects input scripts automatically via Unicode block analysis and routes text to the appropriate processors, inserting transliteration steps when a model trained on one script variant is applied to input in another. Fourth, a lightweight core: mandatory dependencies are minimal (Python 3.9+ and NumPy only); heavy dependencies (PyTorch, Stanza, Hugging Face transformers, hfst) are optional extras installable via pip install turkicnlp[stanza], [nllb], or [all]. Fifth, license isolation: the library core is Apache-2.0 licensed; and other components are downloaded separately at runtime and never bundled with the pip distribution. Sixth, CoNLL-U interoperability: native CoNLL-U import and export ensures full compatibility with the Universal Dependencies ecosystem de Marneffe et al. (2021) and standard downstream tools.
3.2 Pipeline Architecture
The Pipeline class is the central entry point. It accepts a language code, a list of processor names, an optional script override, and per-processor configuration overrides (e.g., morph_backend="apertium"). On initialization the pipeline: (1) determines the working script—auto-detected or user-declared; (2) resolves processor dependencies: each processor class declares REQUIRES annotation layers, so requesting "depparse" automatically adds "tokenize" and "pos" if omitted; (3) consults catalog.json to identify the default backend and model download URL for each processor–language–script combination; and (4) instantiates and loads processors in the following canonical order:
tokenize $\to$ mwt $\to$ morph $\to$ pos $\to$ lemma $\to$ depparse $\to$ ner $\to$ embeddings $\to$ sentiment $\to$ translate
When called on a text string, the pipeline: creates a Document; auto-detects the script if not declared; optionally transliterates the text to the script variant required by the available model; passes the document through each processor sequentially; and optionally back-transliterates output annotations (e.g., lemmas) to the original script. Batch processing is supported via pipeline.batch(texts), and file processing via pipeline.process_file(input, output, format), accepting conllu and json output formats. Figure 2 shows the component structure of the library.
Script ScriptDetector Transliterator Pipeline lang $\cdot$ processors[] script $\cdot$ config Model Mgmt ModelRegistry Downloader catalog.json Processor Chain tokenize $\to$ mwt $\to$ morph $\to$ pos $\to$ lemma $\to$ depparse $\to$ ner $\to$ embeddings $\to$ translate orchestrates Document Model Document $\to$ Sentence $\hookrightarrow$ Token / Word / Span produces
Figure 2: High-level component architecture of TurkicNLP. The Pipeline orchestrates a sequential Processor chain; the Script subsystem handles automatic script detection and transliteration; Model Management resolves and downloads model artifacts via catalog.json; the Document Model provides a hierarchical output representation.
3.3 Document Data Model
TurkicNLP uses a Stanza-compatible hierarchical data model: Document $\to$ Sentence $\to$ Token $\to$ Word $\to$ Span. Each Word carries: text, lemma, upos (Universal POS tag), xpos (language-specific tag), feats (UD morphological features), head and deprel (dependency arc), deps (enhanced dependencies), start_char/end_char, and ner label. Named entity spans are represented as Span objects with label, start_char, and end_char. The Document additionally records script (detected or declared), embedding (document-level vector from the NLLB backend), translation (output from the translation backend), and _processor_log tracing which backends processed the document.
Listing 1 illustrates standard pipeline usage. CoNLL-U output is produced via doc.to_conllu().
Listing 1: Pipeline usage and annotation access (Turkish).
import turkicnlp
turkicnlp.download("tur")
nlp = turkicnlp.Pipeline(
"tur",
processors=["tokenize","pos","lemma","ner","depparse"])
doc = nlp("Halil dün Ankara’ya gitti")
for w in doc.words:
print(w.text, w.upos, w.lemma, w.ner)
# Halil PROPN halil B-PER
# dün NOUN dün O
# Ankara’ya PROPN ankara B-LOC
# gitti VERB git O
print(doc.to_conllu())
3.4 Processor Interface
Each processor is a subclass of an abstract Processor base class that declares what annotation layers it consumes and produces. This declarative dependency model allows the pipeline to resolve processor ordering automatically and to raise an error early if a required layer is missing. The interface is intentionally minimal: processors implement a load method and a process method, and are otherwise self-contained. This design makes it straightforward for third-party contributors to add new languages, new backends, or new processing approaches without modifying the pipeline core.
3.5 Model Registry and Download Manager
Model metadata is maintained in a central JSON manifest (catalog.json) that records, for each language–script–processor combination, a download URL, an integrity checksum, a license identifier, and a quality tier. On first use the library fetches and verifies only the model files needed for the requested processors; subsequent calls use the local cache. Selective download, programmatic discoverability of available languages and processors, and license-aware storage (keeping components with different license terms in separate cache paths) are supported. The manifest-driven approach decouples model management from library code, allowing new models to be registered without releasing a new library version.
4 Script and Orthography Support
4.1 Script Coverage
TurkicNLP supports four script families: Latin (Latn), Cyrillic (Cyrl), Perso-Arabic (Arab), and Old Turkic Runic (Orkh). Several languages use multiple scripts, and a number of countries are undergoing active official script transitions. Full script assignments per language are given in Table 3. Unicode NFC normalization is applied by default; the Turkish dotted/dotless I distinction (İ/I, i/ı) is handled correctly throughout. Available pretrained neural language models per language are surveyed in Table LABEL:tab:llm-models (Appendix).
4.2 Automatic Script Detection
The script detector identifies the dominant script of a text string through Unicode block analysis. Each alphabetic character is classified into one of the four supported script ranges; non-alphabetic characters (digits, punctuation, whitespace) are excluded. A segment-detection function additionally splits mixed-script documents into contiguous single-script runs. Cyrillic/Latin homoglyphs (e.g., Cyrillic a, e, o) are classified by Unicode codepoint rather than visual appearance. Language-aware validation optionally flags a mismatch between the detected script and a declared language code.
4.3 Transliteration
TurkicNLP implements bidirectional transliteration for eight language–script pairs and two one-way conversions (Table 2). Mapping tables follow official national standards where available: the 2021 Kazakh official Latin alphabet, the 1995 Uzbek standard, the Uyghur Latin Yéziqi (ULY). For Ottoman Turkish a simplified academic Latin-to-Arabic mapping is provided; for Old Turkic, one-way Runic $\to$ Latin covers the Orkhon and Yenisei codepoints per standard Turkological conventions.
Language ISO Direction Standard
Kazakh kaz Cyrl $\leftrightarrow$ Latn Official 2021
Uzbek uzb Cyrl $\leftrightarrow$ Latn Official 1995
Azerbaijani aze Cyrl $\leftrightarrow$ Latn Official 1991
Tatar tat Cyrl $\leftrightarrow$ Latn Zamanälif
Turkmen tuk Cyrl $\leftrightarrow$ Latn Official 1993
Karakalpak kaa Cyrl $\leftrightarrow$ Latn Official 2016
Crimean Tatar crh Cyrl $\leftrightarrow$ Latn Latin std.
Uyghur uig Arab $\leftrightarrow$ Latn ULY
Ottoman Turkish ota Latn $\to$ Arab Academic
Old Turkic otk Orkh $\to$ Latn Turkological
Table 2: Transliteration pairs supported by TurkicNLP.
The engine uses a greedy longest-match algorithm: at each position it tries matches of decreasing length (up to four characters), resolving digraphs and trigraphs before single characters. Capitalization is preserved. Script-aware pipeline routing integrates transliteration automatically: when a neural model is trained on one script variant, the pipeline inserts a forward transliteration step before processing and back-transliteration after, making script conversion transparent to the user. Known limitations include: vowel-harmony-conditioned homographs in Kazakh and Kyrgyz (where a Cyrillic sequence may correspond to two distinct Latin strings depending on context); zero-width non-joiner (U+200C) handling in Uyghur Perso-Arabic, where incorrect normalization can break word boundaries; and Perso-Arabic diacritics (harakat) in Ottoman Turkish, which are stripped by the current implementation rather than mapped to Latin equivalents. These edge cases are documented in the repository and targeted for correction in future releases.
Listing 2: Script detection and Kazakh Cyrillic/Latin transliteration.
from turkicnlp.scripts import Script
from turkicnlp.scripts.detector import detect_script
from turkicnlp.scripts.transliterator import Transliterator
# Detect script
cyrl = "Мен Алматыда турамын." # Kazakh Cyrillic
print(detect_script(cyrl)) # Script.CYRILLIC
# Cyrillic -> Latin (2021 official Kazakh alphabet)
t = Transliterator(
"kaz", source=Script.CYRILLIC, target=Script.LATIN)
latin = t.transliterate(cyrl)
print(latin) # Men Almatyda turamyn.
# Reverse: Latin -> Cyrillic
t_back = Transliterator(
"kaz", source=Script.LATIN, target=Script.CYRILLIC)
print(t_back.transliterate(latin))
4.4 Language and Processor Coverage
Table 3 summarises the script and processor availability for all 24 supported languages. Coverage reflects the state of publicly available resources: Turkish and Kazakh receive the full pipeline including NER; five languages have Stanza neural parsing via their UD treebanks; 20 languages have Apertium FST morphological analysis; and all 24 languages receive NLLB-200 sentence embeddings and machine translation.
Language Script
Tokenize
MWT
Morph (FST)
POS Tag
Lemmatize
Dep. Parse
NER
Transliterate
Embed.
Translate
Oghuz
Turkish Latn
Azerbaijani Latn/Cyrl $\times$ $\times$ $\times$ $\times$
Iranian Azerbaijani Arab $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
Turkmen Latn/Cyrl $\sim$ $\times$ $\times$ $\times$ $\times$
Gagauz Latn $\sim$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
Kipchak
Kazakh Cyrl/Latn
Kyrgyz Cyrl $\times$ $\times$
Tatar Cyrl/Latn $\times$ $\times$ $\times$ $\times$
Bashkir Cyrl $\sim$ $\times$ $\times$ $\times$ $\times$ $\times$
Crimean Tatar Latn/Cyrl $\sim$ $\times$ $\times$ $\times$ $\times$
Karakalpak Latn/Cyrl $\sim$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
Nogai Cyrl $\sim$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
Kumyk Cyrl $\sim$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
Karachay-Balkar Cyrl $\sim$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
Karluk
Uzbek Latn/Cyrl $\times$ $\times$ $\times$ $\times$
Uyghur Arab/Latn $\sim$ $\times$
Siberian
Sakha Cyrl $\sim$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
Altai Cyrl $\sim$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
Tuvan Cyrl $\sim$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
Khakas Cyrl $\sim$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
Oghur
Chuvash Cyrl $\sim$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
Historical
Ottoman Turkish Arab $\times$ $\times$ $\times$ $\times$ $\times$
Old Turkish Orkh $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
Arghu
Khalaj Arab $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
Table 3: Script and processor availability for all 24 TurkicNLP languages. Script codes: Latn = Latin, Cyrl = Cyrillic, Arab = Perso-Arabic, Orkh = Old Turkic Runic; primary script listed first for dual-script languages.  = production/stable, $\sim$ = beta/prototype, $\times$ = not available, – = not applicable.
5 Processing Capabilities
The processing pipeline maps each stage to one or more interchangeable backends. For tokenization, a rule-based segmenter handles Latin- and Cyrillic-script input, with a dedicated Arabic-script tokenizer for right-to-left languages; neural Stanza Qi et al. (2020) tokenizers are used for the five UD-treebank languages (Turkish, Kazakh, Kyrgyz, Uyghur, Ottoman Turkish). Morphological analysis draws on Apertium FST binaries Forcada et al. (2011) for 20 languages. POS tagging, lemmatization, and dependency parsing are provided via Stanza neural models for those five languages. Named entity recognition targets Turkish and Kazakh. Sentence embeddings and machine translation use NLLB-200 NLLB Team et al. (2022) for all 24 languages. Table 3 summarises component availability across all 24 languages; runnable code samples for all supported languages are available in the companion repository. [^6] https://github.com/turkic-nlp/turkic-nlp-code-samples The following subsections describe each stage in detail.
5.1 Tokenization and Multi-Word Token Expansion
TurkicNLP provides three tokenization backends. A rule-based tokenizer handles Latin- and Cyrillic-script input via whitespace segmentation augmented by punctuation attachment rules, clitic splitting, and intra-word apostrophe handling (e.g., Turkish possessive forms). A dedicated Arabic-script tokenizer adds right-to-left support, including zero-width non-joiner splitting for Uyghur and Arabic-script punctuation handling. For the five Turkic languages with Universal Dependencies treebanks—Turkish, Kazakh, Kyrgyz, Uyghur, and Ottoman Turkish—a neural tokenizer built on Stanza Qi et al. (2020) provides high-accuracy tokenization trained on UD gold data. A dedicated multi-word token (MWT) processor applies per-language rule tables covering 21 languages, splitting agglutinated forms before downstream morphological analysis.
5.2 Morphological Analysis
Morphological analysis is provided through Apertium finite-state transducer (FST) binaries compiled for 20 Turkic languages, accessed via Python-native HFST bindings without requiring a system-level Apertium installation. For a given token, the FST returns zero or more analyses—each comprising a lemma and a morphological tag string—which a tag-mapping subsystem converts to Universal Dependencies features and UPOS labels following the conventions established by Tyers and Washington (2015). Full UD mapping is implemented for Turkish, Kazakh, and Tatar; partial coverage is available for the remaining 17 languages. Morphological ambiguity is resolved using a combination of FST-internal HFST weights and heuristic scoring based on syntactic position, POS category priority, and lexicon lookups. This remains a known limitation for agglutinative languages where local context strongly determines the correct reading; a lightweight sequence-based disambiguator is planned for a future release. The FST analyzers additionally support generation mode: given a lemma and feature string, the transducer synthesizes the corresponding surface form, useful for data augmentation tasks. Table 4 lists quality levels per language.
Language ISO Branch Quality
Turkish tur Oghuz Production
Kazakh kaz Kipchak Production
Tatar tat Kipchak Production
Azerbaijani aze Oghuz Stable
Kyrgyz kir Kipchak Stable
Uzbek uzb Karluk Stable
Turkmen tuk Oghuz Beta
Bashkir bak Kipchak Beta
Uyghur uig Karluk Beta
Crimean Tatar crh Kipchak Beta
Chuvash chv Oghur Beta
Sakha sah Siberian Prototype
Karakalpak kaa Kipchak Prototype
Nogai nog Kipchak Prototype
Kumyk kum Kipchak Prototype
Karachay-Balkar krc Kipchak Prototype
Altai alt Siberian Prototype
Tuvan tyv Siberian Prototype
Khakas kjh Siberian Prototype
Gagauz gag Oghuz Prototype
Table 4: Apertium FST analyzer quality levels (based on Apertium lexicon size and UD tag-mapping completeness). Production: estimated >>85% token coverage, full UD FEATS mapping. Stable: >>70% coverage, partial UD mapping. Beta: functional with known lexicon gaps. Prototype: limited coverage, Apertium-native tags only.
5.3 POS Tagging, Lemmatization, and Dependency Parsing
For the five Turkic languages with UD treebanks, TurkicNLP provides neural POS tagging, lemmatization, and dependency parsing via Stanza Qi et al. (2020), producing UPOS, XPOS, morphological features, lemma, and CoNLL-U dependency arcs in a single pass using biaffine attention-based models. Turkish offers seven treebank variants: IMST (general-domain default), BOUN, FrameNet, KeNet, ATIS, Penn, and Tourism; Kazakh uses KTB, Kyrgyz the KTMU treebank, Uyghur the UDT treebank, and Ottoman Turkish the BOUN treebank. The active treebank may be specified as a pipeline configuration parameter. For Apertium-backed morphological analysis on any of the 20 supported languages, the same Pipeline interface applies with morph_backend="apertium" and processors=["tokenize","morph"]; the pipeline then populates w.lemma and w.feats from the FST output (see Listing 1 for the analogous Turkish example).
5.4 Named Entity Recognition
NER is available at production quality for Turkish and Kazakh. The Turkish model, integrated via Stanza, is trained on the Starlang annotated corpus and recognizes four entity types (person, organization, location, miscellaneous). The Kazakh model is trained on KazNERD Yeshpanov et al. (2022), a fine-grained corpus of 112,702 annotated tokens spanning 25 entity classes. Named entity spans are stored as structured objects on the document entity list with character offsets and type labels.
5.5 Sentence Embeddings and Machine Translation
TurkicNLP integrates NLLB-200 NLLB Team et al. (2022) [^7] https://huggingface.co/facebook/nllb-200-distilled-600M for cross-lingual sentence embeddings and machine translation across all 24 supported languages, including low-resource varieties absent from alternatives such as LaBSE Feng et al. (2022) and LASER Heffernan et al. (2022). The embeddings processor extracts document-level representations via processors=["embeddings"], producing a doc.embedding vector for cross-lingual similarity, clustering, or classification. The translation processor accepts a translate_tgt_lang NLLB language code and produces doc.translation; beam size is configurable via the num_beams parameter. Both components are available via pip install turkicnlp[nllb].
6 Conclusion
We have presented TurkicNLP, a unified, open-source Python NLP toolkit designed to address the fragmentation and resource scarcity that has long characterised computational work on the Turkic language family. By providing a single, consistent pipeline interface across 24 languages—spanning tokenization, morphological analysis, POS tagging, lemmatization, dependency parsing, named entity recognition, transliteration, sentence embeddings, and machine translation—the library enables researchers and practitioners to work across the breadth of the Turkic family without assembling custom toolchains for each language. Its script-aware architecture handles the orthographic diversity of the family automatically, and its catalog-driven model management keeps the library core lightweight while supporting selective installation of heavier backends. We hope that TurkicNLP lowers the barrier to entry for work on the 200+ million speakers of Turkic languages, particularly the many languages that currently have limited or no digital NLP infrastructure.
7 Future Directions
Near-term development focuses on expanding neural pipeline coverage to additional languages using available pretrained encoders, adding training infrastructure so researchers can contribute custom models, integrating ASR and TTS components to extend the pipeline to spoken language, and conducting a systematic evaluation campaign across all supported languages and processors.