### v4

# TurkicNLP: An NLP Toolkit for Turkic Languages

Abstract
Natural language processing for the Turkic language family, spoken by over 200 million people across Eurasia, remains fragmented, with most languages lacking unified tooling and resources. We present TurkicNLP, an open-source Python library providing a single, consistent NLP pipeline for Turkic languages across four script families: Latin, Cyrillic, Perso-Arabic, and Old Turkic Runic. The library covers tokenization, morphological analysis, part-of-speech tagging, dependency parsing, named entity recognition, bidirectional script transliteration, cross-lingual sentence embeddings, and machine translation through one language-agnostic API. A modular multi-backend architecture integrates rule-based finite-state transducers and neural models transparently, with automatic script detection and routing between script variants. Outputs follow the CoNLL-U standard for full interoperability and extension. Code and documentation are hosted at https://github.com/turkic-nlp/turkicnlp.
TurkicNLP: An NLP Toolkit for Turkic Languages
Sherzod Hakimov Computational Linguistics University of Potsdam firstnamelastname@gmail.com
1 Introduction
1.1 Background
The Turkic languages form one of the world’s largest language families, comprising over 35 recognized languages spoken by approximately 200 million people across a vast geographic region stretching from Turkey and the Balkans to Siberia and northwest China Johanson (2021). These languages are characterized by shared typological features such as agglutinative morphology, vowel harmony, and a subject-object-verb (SOV) word order (as illustrated in Table 1), which distinguish them significantly from many Indo-European languages Jurafsky and Martin (2026). Furthermore, the Turkic family exhibits remarkable script diversity, utilizing Latin, Cyrillic, and Perso-Arabic alphabets, with historical use of Old Turkic Runic scripts. This linguistic and scriptural richness presents unique challenges and opportunities for Natural Language Processing (NLP).
Language (ISO) Sentence Language (ISO) Sentence
Latin script
Turkish (tur) Ali kitabı gördü. Azerbaijani (aze) Ali kitabı gördü.
Turkmen (tuk) Ali kitaby gördi. Gagauz (gag) Ali kitabı gördü.
Crimean Tatar (crh) Ali kitanı kördi. Karakalpak (kaa) Ali kitaptı ko’rdi.
Uzbek (uzb) Ali kitobni ko’rdi.
Cyrillic script
Kazakh (kaz) Aли кітапты көрді. Kyrgyz (kir) Али китепти көрдү.
Tatar (tat) Али китапны күрде. Bashkir (bak) Али китапты күрҙе.
Nogai (nog) Али китапты коьрди. Kumyk (kum) Али китапны гёрдю.
Karachay-Balkar (krc) Али китапны кёрдю. Altai (alt) Али бичикти кӧрди.
Sakha (sah) Али кинигэни көрдө. Tuvan (tyv) Али номну көрген.
Chuvash (chv) Али кӗнекене курнӑ. Khakas (kjh) Али кітапты көрді.
Perso-Arabic script (right-to-left)
Uyghur (uig) ئەلى كىتابنى كۆردى.
Table 1: The sentence “Ali saw the book” across 20 Turkic languages, illustrating the SOV word order (subject Ali — object book with accusative suffix — verb saw at the end) and the three active script families: Latin, Cyrillic, and Perso-Arabic.
Figure 1: Geographic distribution of the Turkic language family (Source: Wikimedia Commons).
1.2 Motivation
Despite the size and diversity of the Turkic language family, NLP tools and resources remain heavily skewed toward Turkish, with most other languages severely underresourced Joshi et al. (2020). This fragmentation means that while some languages like Turkish have access to mature NLP toolkits such as Zemberek Akın and Akın (2007), and others like Kazakh benefit from specialized tools developed by institutions like ISSAI, there is no single, unified library that covers the breadth of the Turkic family with a consistent API. Existing multilingual platforms like Apertium offer rule-based machine translation and morphological analyzers for many Turkic languages Forcada et al. (2011), but often lack a direct, user-friendly Python API for pipeline integration. The agglutinative nature of Turkic languages demands specialized morphological analysis approaches that go beyond the tokenization and part-of-speech tagging paradigms common in NLP for less morphologically complex languages. Moreover, ongoing script transitions in countries like Kazakhstan and Uzbekistan necessitate robust and unified transliteration tooling to bridge linguistic data across different orthographies. The absence of a comprehensive, accessible, and unified NLP toolkit for Turkic languages impedes research, resource creation, and the development of practical applications, particularly for the numerous low-resource languages within the family.
1.3 Contributions
This work makes the following contributions:
  • Unified Pipeline API: A single Pipeline interface covering tokenization, morphological analysis, POS tagging, lemmatization, dependency parsing, named entity recognition, transliteration, sentence embeddings, and machine translation for 24 Turkic languages—with the same calling convention regardless of language or backend.
  • Script-Aware Architecture: Automatic script detection and routing across Latin, Cyrillic, Perso-Arabic, and Old Turkic Runic, with bidirectional transliteration for 8 language–script pairs.
  • Multi-Backend Integration: Rule-based Apertium FST morphological analyzers for 20 languages and Stanza neural models for 10 languages—five with Universal Dependencies treebanks and five with custom-trained models—selectable without API changes. NLLB-200 provides cross-lingual embeddings and translation for all 24 languages.
  • CoNLL-U Interoperability: A Stanza-compatible hierarchical document model with native CoNLL-U import and export, ensuring compatibility with Universal Dependencies tools and treebanks.
  • Open-Source Release: Apache-2.0 licensed, with documentation and code at https://github.com/turkic-nlp/turkicnlp and companion Jupyter notebooks for all supported languages at https://github.com/turkic-nlp/turkic-nlp-code-samples.
2 Related Work
2.1 General-Purpose NLP Toolkits
General-purpose NLP toolkits have reached high maturity for major European languages but offer limited coverage of the Turkic family. Stanza Qi et al. (2020) provides neural pipeline components trained on Universal Dependencies treebanks for 66+ languages, including five Turkic languages: Turkish, Kazakh, Kyrgyz, Uyghur, and Ottoman Turkish. However, it lacks a morphology-first design, does not support script transliteration, and provides no integration with Apertium FSTs for the remaining Turkic languages. Trankit Nguyen et al. (2021) is a transformer-based multilingual toolkit covering 56 languages via pipeline components built on XLM-R; it overlaps with Stanza on UD-treebank Turkic languages but similarly provides no script conversion or FST morphology. spaCy Honnibal et al. (2020) is an industrial-strength library optimized for efficiency across major European languages but offers no dedicated Turkic NLP support beyond standard Unicode handling. NLTK Bird et al. (2009) is a widely-used educational framework with minimal Turkic-specific components. Polyglot Al-Rfou’ et al. (2013) provides distributed word representations and basic NLP components for 100+ languages, including several Turkic varieties, but does not support morphological analysis, script conversion, or unified pipeline integration. UDPipe Straka et al. (2016) offers trainable pipelines for any CoNLL-U treebank but does not bundle Turkic models and lacks transliteration support. UralicNLP Hämäläinen (2019), built on the Giellatekno/Divvun FST infrastructure, provides a comparable morphology-first, script-aware toolkit for Uralic minority languages; its architecture—FST analyzers accessed through a unified Python API—directly parallels TurkicNLP’s design, though it targets an entirely different language family with different morphological and script profiles. calamanCy Miranda (2023) and PyThaiNLP Phatthiyaphaibun et al. (2023) exemplify a complementary paradigm: language-specific pipelines that consolidate disjoint NLP tools for morphologically complex under-resourced languages behind a unified API. TurkicNLP occupies a similar role—consolidating fragmented tools for under-resourced languages—but extends the paradigm to a 24-language family with integrated multi-backend swappability and script-aware transliteration routing as first-class architectural features.
2.2 Turkic-Specific NLP Tools and Pipelines
Several tools address individual Turkic languages, organized here by language.
Turkish: Oflazer and Saraçlar Oflazer and Saraçlar (2018) provide the definitive reference volume on Turkish NLP, surveying 25 years of work on morphological processing, language modeling, deep parsing, machine translation, treebanks, and lexical resources. Zemberek-NLP Akın and Akın (2007) is a comprehensive Java library for Turkish NLP covering morphological analysis, POS tagging, and spell checking; it is restricted to Turkish, does not expose a pip-installable Python API, and has no cross-family support. TurkishDelightNLP Alecakir et al. (2022) is a neural Turkish NLP pipeline offering tokenization, morphological analysis, dependency parsing, and NER with strong performance on Turkish; it targets Turkish exclusively and does not integrate with other Turkic languages or script conversion. VNLP Turker et al. (2023) is an open-source Python NLP library for Turkish providing morphological analysis, named entity recognition, dependency parsing, and sentiment analysis, including quantitative benchmarks on standard Turkish test sets; like TurkishDelightNLP it is Turkish-only and provides no script handling. Istanbul Technical University (ITU) [^1] https://bbf.itu.edu.tr/en/research/research-labs/nlp-lab maintains a complementary Turkish NLP infrastructure including neural pipelines, the IMST and ITU Web treebanks, TRopBank PropBank, and supporting corpora Eryiğit et al. (2023). The Turkish NLP Suite Altinok (2023) is an open-source non-profit collection that includes spaCy-compatible Turkish models, the SentiTurca sentiment corpus, TrGLUE benchmark, InstructTurca instruction-tuning data, and BellaTurca.
Kazakh: The Institute of Smart Systems and Artificial Intelligence (ISSAI) at Nazarbayev University [^2] https://issai.nu.edu.kz/ has developed one of the most comprehensive Turkic-language NLP ecosystems. Text resources include KazNERD Yeshpanov et al. (2022) (112,702 tokens, 25 entity classes), KazQAD Yeshpanov et al. (2024a), and KazParC Yeshpanov et al. (2024c) (371,902 parallel sentences across Kazakh, English, Russian, and Turkish). The speech strand covers KazakhTTS2 Mussakhojayeva et al. (2022) (271 hours TTS), a multilingual Turkic speech corpus spanning 10 languages Mussakhojayeva et al. (2023a), and the Söyle noise-robust ASR system for 11 Turkic languages Mussakhojayeva et al. (2023b).
Kyrgyz: Alekseev and Turatali (2025) survey the state of Kyrgyz NLP, cataloguing available resources and proposing a development roadmap; they note that despite millions of speakers, Kyrgyz remains severely under-resourced. KyrgyzNER Turatali et al. (2025a) provides 10,900 annotated sentences across 27 entity classes for Kyrgyz NER. Turatali et al. (2025b) present a benchmark for evaluating Kyrgyz LLMs covering commonsense reasoning and multilingual understanding tasks. The Cramer Project The Cramer Project (2024) provides dedicated Kyrgyz speech models: AkylAI-STT-small (Whisper fine-tuned for Kyrgyz ASR) and AkylAI-TTS-mini (Matcha-TTS trained on 13 hours of Kyrgyz speech).
Multiple languages: Apertium Forcada et al. (2011) provides rule-based morphological analyzers and machine translation for 20+ Turkic languages via finite-state transducers, representing the most comprehensive freely available resource for Turkic morphological analysis. Its native interface is command-line-based, loading Apertium FST binaries via the hfst Python package without requiring system-level installation. Abdurakhmonova et al. 2022 present an NLP tool for linguistic analysis of Turkic languages focusing on morphological annotation, though it addresses a limited language set without a Python package interface. Universal Dependencies treebanks for Turkic languages were created by dedicated community efforts: the Turkish IMST treebank Sulubacak et al. (2016), the Tatar NMCTT treebank Taguchi et al. (2022), the Uyghur UDT Eli et al. (2016), the Yakut/Sakha treebank Merzhevich and Ferraz Gerardi (2022), and Old Turkish Derin and Harada (2021), with a parallel aligned treebank for Azerbaijani, Kyrgyz, Turkish, and Uzbek Akhundjanova et al. (2025). These contributions remain language-specific and are not integrated into a unified pipeline toolkit. Mirzakhalov et al. (2021) provide the first large-scale MT study across 22 Turkic languages, contributing a parallel corpus of approximately 1.4 million sentences, bilingual baselines for 26 language pairs, multi-domain test sets, and human evaluation scores. Tukeyev et al. (2026) extend this line of work by developing and cleaning parallel corpora for six Turkic languages (Kazakh, Azerbaijani, Kyrgyz, Turkish, Turkmen, and Uzbek) totalling over 3.8 million sentences, and fine-tuning NLLB-200 to achieve substantial gains in BLEU and chrF over the baseline for low-resource pairs.
2.3 Multilingual Models and Benchmarks
Large-scale multilingual models and language-specific pretrained models now cover a growing number of Turkic languages, organized here by language.
Turkish: BERTurk Schweter (2020) provides multiple Turkish BERT variants (cased, uncased, 128k vocabulary). TURNA Uludoğan et al. (2024) provides a 1.1B-parameter Turkish encoder-decoder language model. Large generative LLMs include Kumru-2B Turker et al. (2024) and Trendyol-LLM Trendyol Tech (2024).
Kazakh: KAZ-LLM Yeshpanov et al. (2024b) (8B and 70B parameters, LLaMA 3.1-based) and Sherkala-Chat Koto et al. (2025) are generative models targeting Kazakh.
Azerbaijani: aLLMA Isbarov et al. (2024) is a language model for Azerbaijani.
Uzbek: UzBERT Mansurov and Mansurov (2021), TahrirchiBERT Mamasaidov et al. (2021), and BERTbek Kuriyozov et al. (2024) cover Uzbek.
Kyrgyz: KyrgyzBERT Metinov (2023) is available for Kyrgyz.
Multilingual: Large-scale multilingual models such as mBERT Devlin et al. (2019) and XLM-R Conneau et al. (2020) provide broad cross-lingual representations for Turkic NLP, but suffer from high tokenizer fertility on agglutinative languages and do not constitute processing pipelines. The mGPT multilingual generative model Shliazhko et al. (2024) covers nine Turkic languages (Azerbaijani, Bashkir, Chuvash, Kazakh, Kyrgyz, Tatar, Turkmen, Uzbek, and Sakha). The NLLB-200 model NLLB Team et al. (2022) covers 200 languages including 14 Turkic variants for machine translation.
Evaluation benchmarks: Evaluation benchmarks for Turkic NLP have advanced considerably. TUMLU Isbarov et al. (2025) covers 9 Turkic languages (Azerbaijani, Crimean Tatar, Turkish, Uyghur, Uzbek, Karakalpak, Kyrgyz, Kazakh, and Tatar) with 38K multiple-choice questions; Kardeş-NLU Senel et al. (2024) benchmarks cross-lingual NLU transfer from Turkish into five related languages (Azerbaijani, Kazakh, Kyrgyz, Uzbek, and Uyghur); TR-MMLU Bayram et al. (2025) provides 6,200 multiple-choice questions across 62 subject categories specifically for benchmarking Turkish LLMs; TurkishMMLU Yüksel et al. (2024) offers over 10,000 curriculum-based multiple-choice questions across 9 subjects spanning sciences, mathematics, and culturally specific topics for Turkish; Cetvel Abrek Er et al. (2025) is a unified 23-task evaluation suite assessing language understanding, generation, and cultural competence of LLMs in Turkish; and TurkBench Toraman et al. (2026) provides a broad multi-task benchmark for evaluating Turkish LLMs across diverse NLP tasks. Turatali et al. (2025b) build a dedicated benchmark for Kyrgyz LLMs targeting commonsense reasoning and multilingual transfer. Veitsman and Hartmann 2025 survey the state of NLP for Central Asian Turkic languages (Kazakh, Kyrgyz, Uzbek, Uyghur, Turkmen), documenting persistent resource gaps and the challenges of cross-lingual transfer from higher-resource Turkic languages. Collectively, these benchmarks underscore the need for unified processing infrastructure to support training and evaluation pipelines.
2.4 Speech Technologies for Turkic Languages
Speech processing resources for Turkic languages have expanded significantly through dedicated corpus construction and multilingual modeling efforts. ISSAI has led the development of both ASR and TTS systems spanning the Turkic family: TurkicASR Mussakhojayeva et al. (2023a) trains multilingual ASR models across 10 Turkic languages (Azerbaijani, Bashkir, Chuvash, Kazakh, Kyrgyz, Sakha, Tatar, Turkish, Uyghur, Uzbek), achieving a 56.7% relative CER reduction through cross-lingual transfer compared to monolingual baselines. TurkicTTS Yeshpanov et al. (2023) extends this multilingual paradigm to speech synthesis for a partially overlapping set of 10 languages (Azerbaijani, Bashkir, Kazakh, Kyrgyz, Sakha, Tatar, Turkish, Turkmen, Uyghur, Uzbek)—notably including Turkmen but not Chuvash. Söyle Mussakhojayeva et al. (2023b) provides noise-robust ASR for 11 Turkic languages by fine-tuning Whisper with data augmentation.
Key speech corpora include KSC2 (1,200 hours of Kazakh; Mussakhojayeva et al. 2021a), TSC (218 hours of Turkish; ISSAI 2024), USC (105 hours of Uzbek; Mussakhojayeva et al. 2021b), TatSC (269 hours of Tatar; ISSAI 2023), TurkmenSpeech (251 hours of Turkmen; Mamedov 2024), and KazakhTTS2 (271 hours; Mussakhojayeva et al. 2022). Mozilla Common Voice Ardila et al. (2020) provides community-recorded speech data for several Turkic languages, with Bashkir ($\sim$265 hours), Turkish ($\sim$129 hours), Kyrgyz ($\sim$44 hours), Tatar ($\sim$29 hours), Chuvash ($\sim$12 hours), Sakha ($\sim$6 hours), and Kazakh ($\sim$1 hour). FLEURS Conneau et al. (2022) covers Azerbaijani, Kazakh, Kyrgyz, Turkish, and Uzbek for few-shot speech evaluation. Dedicated Kyrgyz speech models have also emerged: the Cramer Project The Cramer Project (2024) releases AkylAI-STT-small (Whisper fine-tuned for Kyrgyz) and AkylAI-TTS-mini (Matcha-TTS trained on 13 hours of Kyrgyz speech), providing the first openly available language-specific ASR and TTS systems for Kyrgyz.
2.5 Tokenizer Efficiency Across Turkic Languages
(a) Mean fertility (tokens per word).
(b) 95th-percentile tokens per word.
Figure 2: Tokenizer efficiency across 20 Turkic languages and English for four LLM tokenizers: Qwen3.5, LLaMA 3, OLMo 3, and GPT-5.2. Higher fertility indicates less efficient tokenization, reflecting poorer vocabulary coverage for agglutinative languages.
The choice of tokenizer has a profound impact on inference cost and downstream performance for Turkic languages. We evaluate four widely used LLM tokenizers—Qwen3.5-35B, Meta-LLaMA-3-8B, OLMo-3-1025-7B, and GPT-5.2—across 20 Turkic languages using a parallel set of 20 typologically diverse sentences translated into each language (Figure 2).
English exhibits near-optimal fertility of 1.08–2.09 tokens per word across all tokenizers. Turkish, the best-represented Turkic language, reaches 1.95 (Qwen/GPT-5.2) to 3.03 (LLaMA). For under-represented languages, the “tokenizer tax” is severe: Uyghur fertility ranges from 3.38 (GPT-5.2) to 6.85 (OLMo), Sakha from 3.48 (GPT-5.2) to 5.21 (LLaMA), and Bashkir from 2.54 (GPT-5.2) to 5.68 (LLaMA). The 95th-percentile tokens per word—a proxy for worst-case segmentation of complex agglutinated forms—reaches 11.0 for Uyghur and Tatar under LLaMA, compared to 2.0 for English across all tokenizers.
GPT-5.2 and Qwen3.5 consistently achieve the best Turkic coverage, while LLaMA 3 exhibits the highest fertility across all Turkic languages. These findings underscore the need for tokenizer-aware evaluation when deploying multilingual LLMs on Turkic text and motivate continued development of dedicated Turkic tokenization strategies.
2.6 Resource Collections and Community Initiatives
Several community-driven collections catalogue the growing body of Turkic NLP resources. The Turkish NLP resource list Turkish NLP Community (2023) [^3] https://github.com/maidis/nlp-tr is a community-maintained index of Turkish NLP tools, corpora, treebanks, and pre-trained models. The Turkic Interlingua (TIL) project Syzdykov and et al. (2021)  [^4] https://github.com/turkic-interlingua/til-mt provides parallel corpora across 22 Turkic languages together with baseline systems for machine translation, automatic speech recognition, and text-to-speech, enabling systematic multilingual research across the family. The UD-Turkic research group [^5] https://github.com/ud-turkic works on harmonising Universal Dependencies treebanks across Turkic languages, producing the first fully aligned parallel UD treebanks for Azerbaijani, Kyrgyz, Turkish, and Uzbek Akhundjanova et al. (2025). Organisational and workshop infrastructure complements these datasets: SIGTURK, the ACL Special Interest Group on Turkic NLP SIGTURK Organizing Committee (2026) [^6] https://sigturk.github.io, organises dedicated workshop venues that bring together researchers working on Turkic languages and provide an important venue for disseminating resources and evaluation results. The TurkLang conference series [^7] https://turklang.ieees.org/, held annually since 2013, serves as another key venue dedicated to computational processing of Turkic languages, fostering cross-institutional collaboration across the Turkic-speaking world. Cekinel and Çöltekin (2025) provide a recent systematic survey of Turkic language modeling, reviewing available corpora, models, and benchmarks while highlighting persistent challenges around data scarcity, tokenization, and multilingual transfer across the family.
2.7 Turkic NLP Resource Landscape
Table LABEL:tab:resources-full (Appendix) catalogues datasets, NLP components, and toolkit coverage for all 24 languages, grouped by genetic branch. Table LABEL:tab:llm-models (Appendix) surveys all publicly available pretrained neural language models—encoder-only, generative, and multilingual—for each supported language. Together, these tables motivate the design choices described in the subsequent sections.
3 Library Design and Architecture
3.1 Design Goals
TurkicNLP was designed around six core principles. First, a unified API: a single Pipeline entry point operates identically across all 24 supported languages, abstracting backend differences and enabling users to switch languages without code changes. Second, multi-backend: each processor can be fulfilled by interchangeable backends—rule-based, Apertium FST, or Stanza neural—without modifying user-facing code; backend selection is controlled either via catalog.json defaults or explicit constructor parameters. Third, script-awareness: the library detects input scripts automatically via Unicode block analysis and routes text to the appropriate processors, inserting transliteration steps when a model trained on one script variant is applied to input in another. Fourth, a lightweight core: mandatory dependencies are minimal (Python 3.9+ and NumPy only); heavy dependencies (PyTorch, Stanza, Hugging Face transformers, hfst) are optional extras installable via pip install turkicnlp[stanza], [nllb], or [all]. Fifth, license isolation: the library core is Apache-2.0 licensed; and other components are downloaded separately at runtime and never bundled with the pip distribution. Sixth, CoNLL-U interoperability: native CoNLL-U import and export ensures full compatibility with the Universal Dependencies ecosystem de Marneffe et al. (2021) and standard downstream tools.
3.2 Pipeline Architecture
The Pipeline class is the central entry point. It accepts a language code, a list of processor names, an optional script override, and per-processor configuration overrides (e.g., morph_backend="apertium"). On initialization the pipeline: (1) determines the working script—auto-detected or user-declared; (2) resolves processor dependencies: each processor class declares REQUIRES annotation layers, so requesting "depparse" automatically adds "tokenize" and "pos" if omitted; (3) consults catalog.json to identify the default backend and model download URL for each processor–language–script combination; and (4) instantiates and loads processors in the following canonical order:
tokenize $\to$ mwt $\to$ morph $\to$ pos $\to$ lemma $\to$ depparse $\to$ ner $\to$ embeddings $\to$ sentiment $\to$ translate
When called on a text string, the pipeline: creates a Document; auto-detects the script if not declared; optionally transliterates the text to the script variant required by the available model; passes the document through each processor sequentially; and optionally back-transliterates output annotations (e.g., lemmas) to the original script. Batch processing is supported via pipeline.batch(texts), and file processing via pipeline.process_file(input, output, format), accepting conllu and json output formats. Figure 3 shows the component structure of the library.
Script ScriptDetector Transliterator Pipeline lang $\cdot$ processors[] script $\cdot$ config Model Mgmt ModelRegistry Downloader catalog.json Processor Chain tokenize $\to$ mwt $\to$ morph $\to$ pos $\to$ lemma $\to$ depparse $\to$ ner $\to$ embeddings $\to$ translate orchestrates Document Model Document $\to$ Sentence $\hookrightarrow$ Token / Word / Span produces
Figure 3: High-level component architecture of TurkicNLP. The Pipeline orchestrates a sequential Processor chain; the Script subsystem handles automatic script detection and transliteration; Model Management resolves and downloads model artifacts via catalog.json; the Document Model provides a hierarchical output representation.
3.3 Document Data Model
TurkicNLP uses a Stanza-compatible hierarchical data model: Document $\to$ Sentence $\to$ Token $\to$ Word $\to$ Span. Each Word carries: text, lemma, upos (Universal POS tag), xpos (language-specific tag), feats (UD morphological features), head and deprel (dependency arc), deps (enhanced dependencies), start_char/end_char, and ner label. Named entity spans are represented as Span objects with label, start_char, and end_char. The Document additionally records script (detected or declared), embedding (document-level vector from the NLLB backend), translation (output from the translation backend), and _processor_log tracing which backends processed the document.
Listing 1 illustrates standard pipeline usage. CoNLL-U output is produced via doc.to_conllu().
Listing 1: Pipeline usage and annotation access (Turkish).
import turkicnlp
turkicnlp.download("tur")
nlp = turkicnlp.Pipeline(
"tur",
processors=["tokenize","pos","lemma","ner","depparse"])
doc = nlp("Halil dün Ankara’ya gitti")
for w in doc.words:
print(w.text, w.upos, w.lemma, w.ner)
# Halil PROPN halil B-PER
# dün NOUN dün O
# Ankara’ya PROPN ankara B-LOC
# gitti VERB git O
print(doc.to_conllu())
3.4 Processor Interface
Each processor is a subclass of an abstract Processor base class that declares what annotation layers it consumes and produces. This declarative dependency model allows the pipeline to resolve processor ordering automatically and to raise an error early if a required layer is missing. The interface is intentionally minimal: processors implement a load method and a process method, and are otherwise self-contained. This design makes it straightforward for third-party contributors to add new languages, new backends, or new processing approaches without modifying the pipeline core.
3.5 Model Registry and Download Manager
Model metadata is maintained in a central JSON manifest (catalog.json) that records, for each language–script–processor combination, a download URL, an integrity checksum, a license identifier, and a quality tier. On first use the library fetches and verifies only the model files needed for the requested processors; subsequent calls use the local cache. Selective download, programmatic discoverability of available languages and processors, and license-aware storage (keeping components with different license terms in separate cache paths) are supported. The manifest-driven approach decouples model management from library code, allowing new models to be registered without releasing a new library version.
4 Script and Orthography Support
4.1 Script Coverage
TurkicNLP supports four script families: Latin (Latn), Cyrillic (Cyrl), Perso-Arabic (Arab), and Old Turkic Runic (Orkh). Several languages use multiple scripts, and a number of countries are undergoing active official script transitions. Full script assignments per language are given in Table 3. Unicode NFC normalization is applied by default; the Turkish dotted/dotless I distinction (İ/I, i/ı) is handled correctly throughout. Available pretrained neural language models per language are surveyed in Table LABEL:tab:llm-models (Appendix).
4.2 Automatic Script Detection
The script detector identifies the dominant script of a text string through Unicode block analysis. Each alphabetic character is classified into one of the four supported script ranges; non-alphabetic characters (digits, punctuation, whitespace) are excluded. A segment-detection function additionally splits mixed-script documents into contiguous single-script runs. Cyrillic/Latin homoglyphs (e.g., Cyrillic a, e, o) are classified by Unicode codepoint rather than visual appearance. Language-aware validation optionally flags a mismatch between the detected script and a declared language code.
4.3 Transliteration
TurkicNLP implements bidirectional transliteration for eight language–script pairs and two one-way conversions (Table 2). Mapping tables follow official national standards where available: the 2021 Kazakh official Latin alphabet, the 1995 Uzbek standard, the Uyghur Latin Yéziqi (ULY). For Ottoman Turkish a simplified academic Latin-to-Arabic mapping is provided; for Old Turkic, one-way Runic $\to$ Latin covers the Orkhon and Yenisei codepoints per standard Turkological conventions.
Language ISO Direction Standard
Kazakh kaz Cyrl $\leftrightarrow$ Latn Official 2021
Uzbek uzb Cyrl $\leftrightarrow$ Latn Official 1995
Azerbaijani aze Cyrl $\leftrightarrow$ Latn Official 1991
Tatar tat Cyrl $\leftrightarrow$ Latn Zamanälif
Turkmen tuk Cyrl $\leftrightarrow$ Latn Official 1993
Karakalpak kaa Cyrl $\leftrightarrow$ Latn Official 2016
Crimean Tatar crh Cyrl $\leftrightarrow$ Latn Latin std.
Uyghur uig Arab $\leftrightarrow$ Latn ULY
Ottoman Turkish ota Latn $\to$ Arab Academic
Old Turkic otk Orkh $\to$ Latn Turkological
Table 2: Transliteration pairs supported by TurkicNLP.
The engine uses a greedy longest-match algorithm: at each position it tries matches of decreasing length (up to four characters), resolving digraphs and trigraphs before single characters. Capitalization is preserved. Script-aware pipeline routing integrates transliteration automatically: when a neural model is trained on one script variant, the pipeline inserts a forward transliteration step before processing and back-transliteration after, making script conversion transparent to the user. Known limitations include: vowel-harmony-conditioned homographs in Kazakh and Kyrgyz (where a Cyrillic sequence may correspond to two distinct Latin strings depending on context); zero-width non-joiner (U+200C) handling in Uyghur Perso-Arabic, where incorrect normalization can break word boundaries; and Perso-Arabic diacritics (harakat) in Ottoman Turkish, which are stripped by the current implementation rather than mapped to Latin equivalents. These edge cases are documented in the repository and targeted for correction in future releases.
Listing 2: Script detection and Kazakh Cyrillic/Latin transliteration.
from turkicnlp.scripts import Script
from turkicnlp.scripts.detector import detect_script
from turkicnlp.scripts.transliterator import Transliterator
# Detect script
cyrl = "Мен Алматыда турамын." # Kazakh Cyrillic
print(detect_script(cyrl)) # Script.CYRILLIC
# Cyrillic -> Latin (2021 official Kazakh alphabet)
t = Transliterator(
"kaz", source=Script.CYRILLIC, target=Script.LATIN)
latin = t.transliterate(cyrl)
print(latin) # Men Almatyda turamyn.
# Reverse: Latin -> Cyrillic
t_back = Transliterator(
"kaz", source=Script.LATIN, target=Script.CYRILLIC)
print(t_back.transliterate(latin))
4.4 Language and Processor Coverage
Table 3 summarises the script and processor availability for all 24 supported languages. Coverage reflects the state of publicly available resources: Turkish and Kazakh receive the full pipeline including NER; ten languages have Stanza neural parsing via UD treebanks (five official, five custom-trained); 20 languages have Apertium FST morphological analysis; a multilingual Glot500-based neural morphological analyzer and lemmatizer covers 20 languages; and all 24 languages receive NLLB-200 sentence embeddings and machine translation.
Language Script
Tokenize
MWT
Morph (FST)
Morph (Neu)
POS Tag
Lemmatize
Dep. Parse
NER
Transliterate
Embed.
Translate
Oghuz
Turkish Latn
Azerbaijani Latn/Cyrl $\times$
South Azerb. Arab $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
Turkmen Latn/Cyrl $\sim$ $\times$
Gagauz Latn $\sim$ $\sim$ $\sim$ $\sim$ $\times$ $\times$ $\times$ $\times$
Kipchak
Kazakh Cyrl/Latn
Kyrgyz Cyrl $\times$ $\times$
Tatar Cyrl/Latn $\times$
Bashkir Cyrl $\sim$ $\times$ $\times$
Crimean Tatar Latn/Cyrl $\sim$ $\sim$ $\sim$ $\sim$ $\times$ $\times$
Karakalpak Latn/Cyrl $\sim$ $\sim$ $\sim$ $\sim$ $\sim$ $\times$ $\times$ $\times$
Nogai Cyrl $\sim$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
Kumyk Cyrl $\sim$ $\sim$ $\sim$ $\sim$ $\sim$ $\times$ $\times$ $\times$ $\times$
Karachay-Balkar Cyrl $\sim$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
Karluk
Uzbek Latn/Cyrl $\times$
Uyghur Arab/Latn $\sim$ $\times$
Siberian
Sakha Cyrl $\sim$ $\sim$ $\sim$ $\sim$ $\sim$ $\times$ $\times$ $\times$ $\times$
Altai Cyrl $\sim$ $\sim$ $\sim$ $\sim$ $\times$ $\times$ $\times$ $\times$ $\times$
Tuvan Cyrl $\sim$ $\sim$ $\sim$ $\sim$ $\times$ $\times$ $\times$ $\times$ $\times$
Khakas Cyrl $\sim$ $\sim$ $\sim$ $\sim$ $\times$ $\times$ $\times$ $\times$ $\times$
Oghur
Chuvash Cyrl $\sim$ $\sim$ $\sim$ $\sim$ $\times$ $\times$ $\times$ $\times$ $\times$
Historical
Ottoman Turkish Arab $\times$ $\times$ $\times$ $\times$ $\times$
Old Turkish Orkh $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
Arghu
Khalaj Arab $\times$ $\sim$ $\sim$ $\sim$ $\times$ $\times$ $\times$ $\times$ $\times$
Table 3: Script and processor availability for all 24 TurkicNLP languages. Script codes: Latn = Latin, Cyrl = Cyrillic, Arab = Perso-Arabic, Orkh = Old Turkic Runic; primary script listed first for dual-script languages. Morph (Neu) = Glot500-based neural morphological analyzer and lemmatizer. POS Tag, Lemmatize, and Dep. Parse reflect combined Stanza and Glot500 backend coverage.  = production/stable, $\sim$ = beta/prototype/zero-shot, $\times$ = not available, – = not applicable.
5 Processing Capabilities
The processing pipeline maps each stage to one or more interchangeable backends. For tokenization, a rule-based segmenter handles Latin- and Cyrillic-script input, with a dedicated Arabic-script tokenizer for right-to-left languages; neural Stanza Qi et al. (2020) tokenizers are used for the five UD-treebank languages (Turkish, Kazakh, Kyrgyz, Uyghur, Ottoman Turkish) and for five additional languages with custom-trained models (Uzbek, Azerbaijani, Turkmen, Tatar, Bashkir). Morphological analysis draws on Apertium FST binaries Forcada et al. (2011) for 20 languages. POS tagging, lemmatization, and dependency parsing are provided via Stanza neural models for those ten languages. Two multilingual neural models built on a frozen Glot500 Imani et al. (2023) backbone extend neural POS tagging and dependency parsing to 15 languages and morphological analysis with lemmatization to 23 languages. Named entity recognition targets Turkish and Kazakh. Sentence embeddings and machine translation use NLLB-200 NLLB Team et al. (2022) for all 24 languages. Table 3 summarises component availability across all 24 languages; runnable code samples for all supported languages are available in the companion repository. [^8] https://github.com/turkic-nlp/turkic-nlp-code-samples The following subsections describe each stage in detail.
5.1 Tokenization and Multi-Word Token Expansion
TurkicNLP provides three tokenization backends. A rule-based tokenizer handles Latin- and Cyrillic-script input via whitespace segmentation augmented by punctuation attachment rules, clitic splitting, and intra-word apostrophe handling (e.g., Turkish possessive forms). A dedicated Arabic-script tokenizer adds right-to-left support, including zero-width non-joiner splitting for Uyghur and Arabic-script punctuation handling. For the five Turkic languages with Universal Dependencies treebanks—Turkish, Kazakh, Kyrgyz, Uyghur, and Ottoman Turkish—a neural tokenizer built on Stanza Qi et al. (2020) provides high-accuracy tokenization trained on UD gold data. Custom-trained Stanza tokenizers additionally support Uzbek, Azerbaijani, Turkmen, Tatar, and Bashkir. A dedicated multi-word token (MWT) processor applies per-language rule tables covering 21 languages, splitting agglutinated forms before downstream morphological analysis.
5.2 Morphological Analysis
Morphological analysis is provided through Apertium finite-state transducer (FST) binaries compiled for 20 Turkic languages, accessed via Python-native HFST bindings without requiring a system-level Apertium installation. For a given token, the FST returns zero or more analyses—each comprising a lemma and a morphological tag string—which a tag-mapping subsystem converts to Universal Dependencies features and UPOS labels following the conventions established by Tyers and Washington (2015). Full UD mapping is implemented for Turkish, Kazakh, and Tatar; partial coverage is available for the remaining 17 languages. Morphological ambiguity is resolved using a combination of FST-internal HFST weights and heuristic scoring based on syntactic position, POS category priority, and lexicon lookups. This remains a known limitation for agglutinative languages where local context strongly determines the correct reading; a lightweight sequence-based disambiguator is planned for a future release. The FST analyzers additionally support generation mode: given a lemma and feature string, the transducer synthesizes the corresponding surface form, useful for data augmentation tasks. Table 4 lists quality levels per language.
Language ISO Branch Quality
Turkish tur Oghuz Production
Kazakh kaz Kipchak Production
Tatar tat Kipchak Production
Azerbaijani aze Oghuz Stable
Kyrgyz kir Kipchak Stable
Uzbek uzb Karluk Stable
Turkmen tuk Oghuz Beta
Bashkir bak Kipchak Beta
Uyghur uig Karluk Beta
Crimean Tatar crh Kipchak Beta
Chuvash chv Oghur Beta
Sakha sah Siberian Prototype
Karakalpak kaa Kipchak Prototype
Nogai nog Kipchak Prototype
Kumyk kum Kipchak Prototype
Karachay-Balkar krc Kipchak Prototype
Altai alt Siberian Prototype
Tuvan tyv Siberian Prototype
Khakas kjh Siberian Prototype
Gagauz gag Oghuz Prototype
Table 4: Apertium FST analyzer quality levels (based on Apertium lexicon size and UD tag-mapping completeness). Production: estimated >>85% token coverage, full UD FEATS mapping. Stable: >>70% coverage, partial UD mapping. Beta: functional with known lexicon gaps. Prototype: limited coverage, Apertium-native tags only.
5.3 POS Tagging, Lemmatization, and Dependency Parsing
For the five Turkic languages with UD treebanks, TurkicNLP provides neural POS tagging, lemmatization, and dependency parsing via Stanza Qi et al. (2020), producing UPOS, XPOS, morphological features, lemma, and CoNLL-U dependency arcs in a single pass using biaffine attention-based models. Turkish offers seven treebank variants: IMST (general-domain default), BOUN, FrameNet, KeNet, ATIS, Penn, and Tourism; Kazakh uses KTB, Kyrgyz the KTMU treebank, Uyghur the UDT treebank, and Ottoman Turkish the BOUN treebank. The active treebank may be specified as a pipeline configuration parameter. In addition, custom-trained Stanza models provide full neural pipeline support (tokenization, POS tagging, lemmatization, and dependency parsing) for Uzbek, Azerbaijani, Turkmen, Tatar, and Bashkir, extending neural parsing coverage to ten Turkic languages in total. For Apertium-backed morphological analysis on any of the 20 supported languages, the same Pipeline interface applies with morph_backend="apertium" and processors=["tokenize","morph"]; the pipeline then populates w.lemma and w.feats from the FST output (see Listing 1 for the analogous Turkish example).
5.4 Multilingual Neural Models (Glot500)
In addition to the per-language Stanza models, TurkicNLP provides two multilingual neural models that share a common architecture: a frozen Glot500 Imani et al. (2023) encoder serves as the backbone, followed by trainable per-script adapters (Latin, Cyrillic, Arabic), per-language embedding vectors, and a shared two-layer BiLSTM, which feeds into task-specific heads. This design forces cross-lingual transfer through shared weights while accommodating script diversity through lightweight adapter modules.
POS Tagger & Dependency Parser.
The first model jointly predicts UPOS tags and labeled dependency arcs. It was trained in Phase 1 (frozen backbone with only adapters, language embeddings, BiLSTM, and task heads trained) on all available Universal Dependencies treebanks for 10 Turkic languages: Turkish (9 treebanks, 82K sentences), Azerbaijani, Turkmen, Bashkir, Uyghur, Kyrgyz, Ottoman Turkish, Tatar, Uzbek, and Kazakh—totaling 116K sentences from 20 UD treebanks. The dependency parsing head uses biaffine attention Dozat and Manning (2017). Five additional languages—Karakalpak, Kumyk, Sakha, Karachay-Balkar, and Nogai—are supported in zero-shot mode via proxy language embeddings from closely related trained languages.
Morphological Analyzer & Lemmatizer.
The second model predicts UPOS tags, 69 UD morphological feature–value pairs (as a multi-label sigmoid classification), and lemmas via edit-script classification (a character-level CNN that predicts surface-to-lemma transformations). Training proceeded in two phases: Phase 1 with the backbone frozen and Phase 2 with the top encoder layers partially unfrozen using discriminative learning rates. The training data combines three complementary sources: (1) UD treebanks for 10 languages providing words in sentence context with gold annotations; (2) morphological paradigm tables from UniMorph Sylak-Glassman (2016) covering 13 Turkic languages with approximately 1.9M entries; and (3) inflectional paradigms extracted from Wiktionary using Wiktextract Ylonen (2022) for 20 languages with approximately 4.8M form triples. This combination extends coverage to 20 Turkic languages in total: 10 with full UD training data and 10 additional languages (Crimean Tatar, Sakha, Tuvan, Khakas, Altai, Kumyk, Gagauz, Chuvash, Khalaj, and Northern Altai) trained on paradigm data only. Three additional languages—Karakalpak, Karachay-Balkar, and Nogai—are supported in zero-shot mode, bringing the total to 23 languages.
Both models use temperature-based sampling (T=2.0T\!=\!2.0) during training to balance the heavily skewed data distribution (Turkish: 82K vs. Kazakh: 1K sentences) and are available via pip install turkicnlp[transformers].
5.5 Named Entity Recognition
NER is available at production quality for Turkish and Kazakh. The Turkish model, integrated via Stanza, is trained on the Starlang annotated corpus and recognizes four entity types (person, organization, location, miscellaneous). The Kazakh model is trained on KazNERD Yeshpanov et al. (2022), a fine-grained corpus of 112,702 annotated tokens spanning 25 entity classes. Named entity spans are stored as structured objects on the document entity list with character offsets and type labels.
5.6 Sentence Embeddings and Machine Translation
TurkicNLP integrates NLLB-200 NLLB Team et al. (2022) [^9] https://huggingface.co/facebook/nllb-200-distilled-600M for cross-lingual sentence embeddings and machine translation across all 24 supported languages, including low-resource varieties absent from alternatives such as LaBSE Feng et al. (2022) and LASER Heffernan et al. (2022). The embeddings processor extracts document-level representations via processors=["embeddings"], producing a doc.embedding vector for cross-lingual similarity, clustering, or classification. The translation processor accepts a translate_tgt_lang NLLB language code and produces doc.translation; beam size is configurable via the num_beams parameter. Both components are available via pip install turkicnlp[nllb].
6 Evaluation
6.1 Morphological Analysis
To assess the two morphological analysis backends—Apertium FST (rule-based) and the Glot500-based neural analyzer—we designed a set of 20 parallel challenge sentences targeting core grammatical phenomena: case marking (genitive, accusative, locative, ablative, dative), verbal categories (past, present progressive, future, negation, imperative, conditional, evidential, passive), and additional features (plural nouns, possessives, pronouns, reciprocals, numerals, converbs). Each sentence was translated into 20 Turkic languages (Appendix A). For each sentence, we check whether the expected UD feature (e.g., Case=Acc, Tense=Past) appears in at least one analysed token.
Table 5 reports three metrics per backend and language: Det% (detection accuracy)—the fraction of 20 challenge sentences where the expected grammatical phenomenon was correctly identified; Ana% (analysis coverage)—the fraction of content words that received a UPOS tag other than X; and Feat% (feature coverage)—the fraction of content words annotated with at least one morphological feature. Lemma coverage is omitted from the table as it is 100% for all languages and backends except Uyghur neural (65%).
Apertium FST Neural (Glot500)
Language ISO Scr Det% Ana% Feat% Det% Ana% Feat%
Turkish tur Latn 75 98 83 80 100 90
Azerbaijani aze Latn 50 53 42 75 100 79
Turkmen tuk Latn 70 69 49 75 100 90
Gagauz gag Latn 70 89 73 75 100 94
Kazakh kaz Cyrl 60 97 69 70 100 82
Kyrgyz kir Cyrl 60 100 76 60 100 92
Tatar tat Cyrl 50 89 45 75 100 89
Bashkir bak Cyrl 60 97 56 75 100 89
Crimean Tatar crh Latn 65 95 70 75 100 89
Karakalpak kaa Latn 45 52 32 50 100 86
Nogai nog Cyrl 50 72 56 55 100 93
Kumyk kum Cyrl 55 84 69 70 100 95
Karachay-Balkar krc Cyrl 10 3 3 70 100 97
Uzbek uzb Latn 75 98 80 55 100 87
Uyghur uig Arab 60 100 63 55 100 60
Sakha sah Cyrl 45 73 63 60 100 85
Altai alt Cyrl 30 36 25 75 100 89
Tuvan tyv Cyrl 55 92 69 75 100 89
Khakas kjh Cyrl 10 6 3 85 100 89
Chuvash chv Cyrl 70 100 57 60 100 92
Average 53 75 54 68 100 88
Table 5: Morphological analysis evaluation on 20 challenge sentences per language (Appendix A). Det%: fraction of sentences where the target UD feature was detected. Ana%: fraction of content words receiving a valid UPOS tag. Feat%: fraction of content words with morphological features. Bold marks the better backend per language and metric. Languages are grouped by branch: Oghuz, Kipchak, Karluk, Siberian, and Oghur. The neural backend wins on 17/20 languages (by detection accuracy); Apertium leads for Uzbek, Uyghur, and Chuvash.
The neural backend achieves higher average detection accuracy (68% vs. 53%) and substantially better feature coverage (88% vs. 54%), benefiting from its training on UD treebanks, UniMorph paradigms, and Wiktionary data. It wins on 17 of 20 languages. The Apertium FST backend outperforms on three languages: Uzbek and Uyghur (where Apertium has mature transducers but the neural model lacks sufficient training data) and Chuvash (where the Oghur branch’s divergent morphology challenges the neural model’s cross-lingual transfer). The most dramatic gaps appear for prototype-level Apertium languages—Khakas (10% vs. 85% detection) and Karachay-Balkar (10% vs. 70%)—where the FST lexicons remain small. Conversely, production-level Apertium languages (Turkish: 75%, Uzbek: 75%) approach neural performance on detection, indicating that the two backends are complementary: rule-based transducers excel where hand-curated lexicons exist, while neural transfer fills gaps for under-resourced languages.
6.2 POS Tagging and Dependency Parsing
We evaluate POS tagging and dependency parsing by comparing the per-language Stanza models (§5.3) against the multilingual Glot500 joint POS tagger and dependency parser (§5.4). Test data for Turkish, Kazakh, Kyrgyz, Uyghur, Uzbek, Karakalpak, Kumyk, and Sakha are drawn from the Parallel Universal Dependencies Treebanks for Turkic Languages Akhundjanova et al. (2025); [^10] https://github.com/ud-turkic/parallel test sets for Azerbaijani, Bashkir, Turkmen, and Tatar are sourced from a complementary generated UD collection. [^11] https://github.com/turkic-nlp/generated-ud-data All test files use gold tokenization for the parsing evaluation; tokenization F1 is reported separately for the Stanza backend.
Table 6 reports UPOS accuracy, UAS, and LAS for both backends across 12 languages. The Stanza backend provides per-language models for 9 languages; the Glot500 multilingual model covers all 12, including three zero-shot languages (Karakalpak via Uzbek, Kumyk via Tatar, Sakha via Bashkir proxy embeddings).
Stanza (per-language) Glot500 (multilingual)
Language ISO $n$ UPOS UAS LAS UPOS UAS LAS
Turkish tur 877 85.1 75.3 63.5 86.9 84.9 70.1
Azerbaijani aze 869 84.9 72.3 60.8 87.0 76.2 64.6
Turkmen tuk 825 86.1 76.8 67.8 85.8 77.8 68.4
Kazakh kaz 851 87.9 67.4 57.4 88.8 71.5 59.7
Kyrgyz kir 1166 71.1 50.2 29.4 79.0 63.8 40.9
Bashkir bak 814 87.0 72.8 62.5 89.6 76.1 64.1
Tatar tat 29 86.2 69.2 57.7 79.3 77.8 70.4
Uyghur uig 758 76.9 77.1 62.3 79.7 81.7 66.9
Uzbek uzb 905 71.8 67.2 46.3 87.6 81.4 68.1
Karakalpak kaa 1126 82.1 67.5 51.0
Kumyk kum 144 78.5 60.7 45.9
Sakha sah 16 93.8 60.0 46.7
Avg (9 shared) 81.9 69.8 56.4 84.9 76.8 63.6
Table 6: POS tagging and dependency parsing evaluation on Parallel Turkic UD test sets Akhundjanova et al. (2025). $n$: number of tokens in the test set. UPOS: POS tagging accuracy. UAS/LAS: unlabeled/labeled attachment scores. Bold marks the better backend per language and metric. Zero-shot languages (shaded): no per-language Stanza model available; Glot500 uses proxy language embeddings. Averages are computed over the 9 languages where both backends are available.
The Glot500 multilingual model outperforms the per-language Stanza models on average across all three metrics: UPOS accuracy (84.9% vs. 81.9%), UAS (76.8% vs. 69.8%), and LAS (63.6% vs. 56.4%). The multilingual model wins on 7 of 9 shared languages for LAS, with particularly large gains for Uzbek (+21.8 LAS), Kyrgyz (+11.5), and Tatar (+12.7). Stanza retains a narrow POS advantage on Turkmen (86.1 vs. 85.8) and Tatar (86.2 vs. 79.3), although the Tatar test set contains only 29 tokens. The three zero-shot languages demonstrate that cross-lingual transfer via proxy embeddings provides usable parsing even without any target-language training data: Karakalpak achieves 82.1% UPOS and 51.0 LAS via Uzbek, while Kumyk (via Tatar) and Sakha (via Bashkir) show lower but functional parsing quality. Tokenization F1 for the Stanza backend exceeds 96% on all tested languages, reaching 100% on Tatar and 99.9% on Uyghur.
Systematic evaluation of NER (entity-level F1 for Turkish and Kazakh) is planned for a future version of this paper.
7 Conclusion
We have presented TurkicNLP, a unified, open-source Python NLP toolkit designed to address the fragmentation and resource scarcity that has long characterised computational work on the Turkic language family. By providing a single, consistent pipeline interface across 24 languages—spanning tokenization, morphological analysis, POS tagging, lemmatization, dependency parsing, named entity recognition, transliteration, sentence embeddings, and machine translation—the library enables researchers and practitioners to work across the breadth of the Turkic family without assembling custom toolchains for each language. Its script-aware architecture handles the orthographic diversity of the family automatically, and its catalog-driven model management keeps the library core lightweight while supporting selective installation of heavier backends. We hope that TurkicNLP lowers the barrier to entry for work on the 200+ million speakers of Turkic languages, particularly the many languages that currently have limited or no digital NLP infrastructure.
8 Future Directions
Near-term development focuses on expanding neural pipeline coverage to additional languages using available pretrained encoders, adding training infrastructure so researchers can contribute custom models, integrating ASR and TTS components to extend the pipeline to spoken language, and conducting a systematic evaluation campaign across all supported languages and processors.
Appendix A Morphological Challenge Sentences
Tables 918 present a parallel set of 20 morphological challenge sentences across English and 20 Turkic languages. Each sentence exercises a specific grammatical phenomenon central to Turkic morphology. These sentences serve as diagnostic inputs for the morphological analysis pipeline described in §5.
Genitive-Possessive gen_psor
eng Ali’s book is on the table.
alt Алинин бичиги столдо.
aze Ali’nin kitabı masadadır.
bak Алиҙың китабы өҫтәлдә.
chv Алин кӗнеки сӗтел ҫинче.
crh Ali’niñ kitabı masada.
gag Ali’nin kitabı masada.
kaa Ali’nin’ kitabı stolda.
kaz Әлидің кітабы үстелде.
kir Алинин китеби үстөлдө.
kjh Алиның кітабы столда.
krc Али́ни китабы́ столдады.
kum Алини китабы столда.
nog Алидинъ китабы столда.
sah Али кинигэтэ остуолга.
tat Алиның китабы өстәлдә.
tuk Aliniň kitaby stoluň üstünde.
tur Ali’nin kitabı masada.
tyv Алинин ному столда.
uig ئەلىنىڭ كىتابى ئۈستەلدە.
uzb Alining kitobi stolda.
Accusative Object acc_object
eng Ali saw the book.
alt Али бичикти кӧрди.
aze Ali kitabı gördü.
bak Али китапты күрҙе.
chv Али кӗнекене курнӑ.
crh Ali kitanı kördi.
gag Ali kitabı gördü.
kaa Ali kitaptı ko’rdi.
kaz Әли кітапты көрді.
kir Али китепти көрдү.
kjh Али кітапты көрді.
krc Али́ китапны́ кёрдю́.
kum Али китапны гёрдю.
nog Али китапты коьрди.
sah Али кинигэни көрдө.
tat Али китапны күрде.
tuk Ali kitaby gördi.
tur Ali kitabı gördü.
tyv Али номну көрген.
uig ئەلى كىتابنى كۆردى.
uzb Ali kitobni ko’rdi.
Table 9: Challenge sentences: Genitive-Possessive and Accusative Object.
Locative locative
eng Ali is waiting at school.
alt Али школдо сакып јат.
aze Ali məktəbdə gözləyir.
bak Али мәктәптә көтә.
chv Али шкулта кӗтет.
crh Ali mektepte bekley.
gag Ali şkolada bekler.
kaa Ali mektepte ku’tip otir.
kaz Әли мектепте күтіп тұр.
kir Али мектепте күтүп жатат.
kjh Али школда күтіп тұр.
krc Али́ шко́лда сакъла́йды.
kum Али школда гёзлеп тура.
nog Али мектепте куьтеди.
sah Али оскуолаҕа кэтэһэр.
tat Али мәктәптә көтә.
tuk Ali mekdepde garaşýar.
tur Ali okulda bekliyor.
tyv Али школада манап олур.
uig ئەلى مەكتەپتە ساقلاۋاتىدۇ.
uzb Ali maktabda kutyapti.
Ablative ablative
eng Ali left the school.
alt Али школдоҥ чыкты.
aze Ali məktəbdən çıxdı.
bak Али мәктәптән сыҡты.
chv Али шкултан тухнӑ.
crh Ali mektepten çıqtı.
gag Ali şkoladan çıktı.
kaa Ali mektepten shiqti.
kaz Әли мектептен шықты.
kir Али мектептен чыкты.
kjh Али школдан шықты.
krc Али́ шко́лдан чы́кты.
kum Али школдан чыкъды.
nog Али мектептен шыкты.
sah Али оскуолаттан таҕыста.
tat Али мәктәптән чыкты.
tuk Ali mekdepden çykdy.
tur Ali okuldan çıktı.
tyv Али школадаан үнген.
uig ئەلى مەكتەپتىن چىقتى.
uzb Ali maktabdan chiqdi.
Table 10: Challenge sentences: Locative and Ablative.
Dative dative
eng Ali entered the school.
alt Али школго кирди.
aze Ali məktəbə girdi.
bak Али мәктәпкә инде.
chv Али шкула кӗнӗ.
crh Ali mektepke kirdi.
gag Ali şkolaya girdi.
kaa Ali mektepke kirdi.
kaz Әли мектепке кірді.
kir Али мектепке кирди.
kjh Али школға кірді.
krc Али́ шко́лгъа кирди́.
kum Али школгъа гирди.
nog Али мектепке кирди.
sah Али оскуолаҕа киирдэ.
tat Али мәктәпкә керде.
tuk Ali mekdebe girdi.
tur Ali okula girdi.
tyv Али школаже кирген.
uig ئەلى مەكتەپكە كىردى.
uzb Ali maktabga kirdi.
Plural Noun plural_noun
eng The children played in the park.
alt Балдар паркта ойнодылар.
aze Uşaqlar parkda oynadılar.
bak Балалар паркта уйнанылар.
chv Ачасем паркра вылянӑ.
crh Balalar parkta oynadılar.
gag Uşaklar parkta oynadılar.
kaa Balalar parkta oynadi.
kaz Балалар саябақта ойнады.
kir Балдар паркта ойношту.
kjh Балалар паркта ойнадылар.
krc Сабийле́ па́ркда ойна́дыла.
kum Яшлар паркта ойнадылар.
nog Балалар паркта ойнадылар.
sah Оҕолор пааркаҕа оонньоотулар.
tat Балалар паркта уйнадылар.
tuk Çagalar parkda oýnadylar.
tur Çocuklar parkta oynadı.
tyv Уруглар паркта ойнааннар.
uig بالىلار باغچىدا ئوينىدى.
uzb Bolalar parkda o’ynashdi.
Table 11: Challenge sentences: Dative and Plural Noun.
1st Person Possessive psor_1sg
eng My house is big.
alt Мениҥ айылым јаан.
aze Mənim evim böyükdür.
bak Минең өйөм ҙур.
chv Манӑн кил пысӑк.
crh Meniñ evim büyük.
gag Benim evim büük.
kaa Menin’ u’yim u’lken.
kaz Менің үйім үлкен.
kir Менин үйүм чоң.
kjh Менің үйім үлкен.
krc Мени́ юйю́м уллу́ду.
kum Мени уьюм уллу.
nog Меним уьйим уьйкен.
sah Мин дьиэм улахан.
tat Минем өем зур.
tuk Meniň öýüm uly.
tur Benim evim büyük.
tyv Мээң бажыңым улуг.
uig مېنىڭ ئۆيۈم چوڭ.
uzb Mening uyim katta.
Past Tense past_tense
eng Ali came yesterday.
alt Али тӱӱне келди.
aze Ali dünən gəldi.
bak Али кисә килде.
chv Али ӗнер килнӗ.
crh Ali tünevin keldi.
gag Ali dün geldi.
kaa Ali keshe keldi.
kaz Әли кеше келді.
kir Али кечээ келди.
kjh Али кеше келді.
krc Али́ тюне́н келди́.
kum Али тюнегюн гелди.
nog Али кеше келди.
sah Али бэҕэһээ кэллэ.
tat Али кичә килде.
tuk Ali düýn geldi.
tur Ali dün geldi.
tyv Али кежээ келген.
uig ئەلى تۈنۈگۈن كەلدى.
uzb Ali kecha keldi.
Table 12: Challenge sentences: 1st Person Possessive and Past Tense.
Present Progressive present_prog
eng Ali is coming now.
alt Али эмди келип јат.
aze Ali indi gəlir.
bak Али хәҙер килә.
chv Али халӗ килет.
crh Ali şimdi kele.
gag Ali şimdi geler.
kaa Ali ha’zir kelip atir.
kaz Әли қазір келе жатыр.
kir Али азыр келе жатат.
kjh Али қазір келе жатыр.
krc Али́ бусагъатда́ келе́ди.
kum Али гьали геле.
nog Али аьзир келеди.
sah Али билигин кэлэр.
tat Али хәзер килә.
tuk Ali häzir gelýär.
tur Ali şimdi geliyor.
tyv Али ам кел чыр.
uig ئەلى ھازىر كېلىۋاتىدۇ.
uzb Ali hozir kelyapti.
Future Tense future_tense
eng Ali will come tomorrow.
alt Али эртен келер.
aze Ali sabah gələcək.
bak Али иртәгә киләсәк.
chv Али ыран килӗ.
crh Ali yarın kelecek.
gag Ali yarın gelecek.
kaa Ali erteń keledi.
kaz Әли ертең келеді.
kir Али эртең келет.
kjh Али ертең келеді.
krc Али́ тамбла́ келли́кди.
kum Али танга гележек.
nog Али эртен келеек.
sah Али сарсын кэлиэҕэ.
tat Али иртәгә киләчәк.
tuk Ali ertir geler.
tur Ali yarın gelecek.
tyv Али даарта кээр.
uig ئەلى ئەتە كېلىدۇ.
uzb Ali ertaga keladi.
Table 13: Challenge sentences: Present Progressive and Future Tense.
Negation negation
eng Ali did not go to school.
alt Али школго барган јок.
aze Ali məktəbə getmədi.
bak Али мәктәпкә барманы.
chv Али шкула кайман.
crh Ali mektepke barmadı.
gag Ali şkolaya gitmedi.
kaa Ali mektepke barmadi.
kaz Әли мектепке бармады.
kir Али мектепке барган жок.
kjh Али школға бармады.
krc Али́ шко́лгъа барма́ды.
kum Али школгъа бармады.
nog Али мектепке бармады.
sah Али оскуолаҕа барбата.
tat Али мәктәпкә бармады.
tuk Ali mekdebe gitmedi.
tur Ali okula gitmedi.
tyv Али школаже барбаан.
uig ئەلى مەكتەپكە بارمىدى.
uzb Ali maktabga bormadi.
Question question
eng Did Ali go to school?
alt Али школго барды ба?
aze Ali məktəbə getdi mi?
bak Али мәктәпкә барҙымы?
chv Али шкула кайнӑ-и?
crh Ali mektepke bardı mı?
gag Ali şkolaya gitti mi?
kaa Ali mektepke bardı ma?
kaz Әли мектепке барды ма?
kir Али мектепке бардыбы?
kjh Али школға барды ма?
krc Али́ шко́лгъа барды́мы?
kum Али школгъа бардымы?
nog Али мектепке бардыма?
sah Али оскуолаҕа барда дуо?
tat Али мәктәпкә бардымы?
tuk Ali mekdebe gitdimi?
tur Ali okula gitti mi?
tyv Али школаже барды бе?
uig ئەلى مەكتەپكە باردىمۇ؟
uzb Ali maktabga bordimi?
Table 14: Challenge sentences: Negation and Question.
Imperative imperative
eng Ali, open the door!
alt Али, эжикти ач!
aze Ali, qapını aç!
bak Али, ишекте ас!
chv Али, алӑка уҫ!
crh Ali, qapını aç!
gag Ali, kapıyı aç!
kaa Ali, esikti ash!
kaz Әли, есікті аш!
kir Али, эшикти ач!
kjh Али, есікті аш!
krc Али́, эшикни́ ач!
kum Али, эшикни ач!
nog Али, эсикти аш!
sah Али, ааны ас!
tat Али, ишекне ач!
tuk Ali, gapyny aç!
tur Ali, kapıyı aç!
tyv Али, эжикти ажыт!
uig ئەلى، ئىشىكنى ئاچ!
uzb Ali, eshikni och!
Conditional conditional
eng If Ali came, it would be good.
alt Али келзе, јараар эди.
aze Əgər Ali gəlsəydi, yaxşı olardı.
bak Әгәр Али килһә, яҡшы булыр ине.
chv Енчен Али килсен, аван пулнӑ пулӗччӗ.
crh Eger Ali kelse, yahşı olur edi.
gag Eer Ali gelseydi, ii olaceydı.
kaa Eger Ali kelse, jaqsi bolar edi.
kaz Егер Әли келсе, жақсы болар еді.
kir Эгер Али келсе, жакшы болмок.
kjh Егер Али келсе, жақсы болар еді.
krc Али́ келсе́, иги́ боллу́къ эди.
kum Али гелсе, яхшы болар эди.
nog Эгер Али келсе, ийги болар эди.
sah Али кэлбит буолсай, үчүгэй буолуо этэ.
tat Әгәр Али килсә, яхшы булыр иде.
tuk Eger Ali gelse, gowy bolardy.
tur Ali gelse iyi olurdu.
tyv Али келзе, эки боор ийик.
uig ئەگەر ئەلى كەلسە، ياخشى بولاتتى.
uzb Agar Ali kelsa, yaxshi bo’lardi.
Table 15: Challenge sentences: Imperative and Conditional.
Evidential evidential
eng Ali apparently came.
alt Али келип калган.
aze Deyəsən, Ali gəlib.
bak Күрәһең, Али килгән.
chv Али килнӗ курӑнать.
crh Ali kelgen kibi.
gag Ali gelmiş.
kaa Ali kelgen sıyaqlı.
kaz Әли келіпті.
kir Али келиптир.
kjh Әли келіпті.
krc Али́ келгенди́.
kum Али гелгендир.
nog Али келген.
sah Али кэлбит.
tat Али килгән, ахрысы.
tuk Ali gelipdir.
tur Ali gelmiş.
tyv Али келген-дир.
uig ئەلى كېلىپتۇ.
uzb Ali kelibdi.
Passive passive
eng The door was opened.
alt Эжик ачылды.
aze Qapı açıldı.
bak Ишек асылды.
chv Алӑк уҫӑлнӑ.
crh Qapı açıldı.
gag Kapı açıldı.
kaa Esik ashildi.
kaz Есік ашылды.
kir Эшик ачылды.
kjh Есік ашылды.
krc Эшик ачы́лды.
kum Эшик ачылды.
nog Эсик ашылды.
sah Аан аһылынна.
tat Ишек ачылды.
tuk Gapy açyldy.
tur Kapı açıldı.
tyv Эжик ажыттынган.
uig ئىشىك ئېچىلدى.
uzb Eshik ochildi.
Table 16: Challenge sentences: Evidential and Passive.
Reciprocal reciprocal
eng They saw each other.
alt Олор бой-бойын кӧрдилӧр.
aze Onlar bir-birini gördülər.
bak Улар бер-береһен күрҙеләр.
chv Вӗсем пӗр-пӗрне курнӑ.
crh Olar bir-birini kördiler.
gag Onnar biri-birini gördülär.
kaa Olar bir-birin ko’rdi.
kaz Олар бір-бірін көрді.
kir Алар бири-бирин көрүштү.
kjh Олар бір-бірін көрді.
krc Ала́ бир-бирлерин кёрдюле́.
kum Олар бир-бирин гёрдюлер.
nog Олар бир-бирин коьрдилер.
sah Кинилэр бэйэ-бэйэлэрин көрүстүлэр.
tat Алар бер-берсен күрделәр.
tuk Olar biri-birini gördüler.
tur Onlar birbirini gördü.
tyv Олар бот-боттарын көргеннер.
uig ئۇلار بىر-بىرىنى كۆردى.
uzb Ular bir-birlarini ko’rdilar.
Pronouns pronouns
eng We know you.
alt Бис сени билерис.
aze Biz səni tanıyırıq.
bak Беҙ һине беләбеҙ.
chv Эпир сана пӗлетпӗр.
crh Biz seni tanıyırmız.
gag Biz seni tanıyırız.
kaa Biz seni taniymiz.
kaz Біз сені танимыз.
kir Биз сени тааныйбыз.
kjh Біз сені танимыз.
krc Биз сени́ таныйбыз.
kum Биз сени таныйбыз.
nog Биз сени таниймыз.
sah Биһиги эйигин билэбит.
tat Без сине беләбез.
tuk Biz seni tanaýarys.
tur Biz seni tanıyoruz.
tyv Бис сени таныыр бис.
uig بىز سېنى تونۇيمىز.
uzb Biz seni taniymiz.
Table 17: Challenge sentences: Reciprocal and Pronouns.
Numeral numeral
eng Three students came.
alt Ӱч студент келди.
aze Üç tələbə gəldi.
bak Өс студент килде.
chv Виҫӗ студент килнӗ.
crh Üç student keldi.
gag Üç öğrenci geldi.
kaa U’sh student keldi.
kaz Үш студент келді.
kir Үч студент келди.
kjh Үш студент келді.
krc Юч студент келди́.
kum Уьч студент гелди.
nog Уьш студент келди.
sah Үс үөрэнээччи кэллэ.
tat Өч студент килде.
tuk Üç talyp geldi.
tur Üç öğrenci geldi.
tyv Үш сургуул келген.
uig ئۈچ ئوقۇغۇچى كەلدى.
uzb Uchta talaba keldi.
Converb converb
eng Ali spoke while laughing.
alt Али кӱлӱп куучындады.
aze Ali gülərək danışdı.
bak Али көлөп һөйләне.
chv Али кулса калаҫнӑ.
crh Ali külerek laf etti.
gag Ali gülerek laf etti.
kaa Ali ku’lip so’yledi.
kaz Әли күліп сөйледі.
kir Али күлүп сүйлөдү.
kjh Али күліп сөйледі.
krc Али́ кюле́ туруп сёлешди́.
kum Али кюлеп сёйледи.
nog Али куьлип соьйледи.
sah Али күлэн саҥарда.
tat Али көлеп сөйләде.
tuk Ali ýylgyryp gürledi.
tur Ali gülerek konuştu.
tyv Али хүлүмзүрүп чугаалаан.
uig ئەلى كۈلۈپ سۆزلىدى.
uzb Ali kulib gapirdi.
Table 18: Challenge sentences: Numeral and Converb.